OpenStudy (briensmarandache):

cos(cos^-1(x) + sin^-1(x)) rewrite as a algebraic expression

OpenStudy (anonymous):

\[\cos (\cos ^{-1}x+\sin ^{-1}x)\]=\(\cos(\frac{ 1 }{ \cos x }+\frac{ 1 }{ \sin x })\) =\(\cos (\frac{ \sin x + \cos x }{ \cos x \sin x })\) can you do the remaining part?

OpenStudy (briensmarandache):

So I'm not supposed to use the cos x+y identity

OpenStudy (briensmarandache):

So is that the same as 1+ cotx?

OpenStudy (briensmarandache):

@kryton1212

OpenStudy (anonymous):

oh umm... i haven't learnt cotx yet. i am used to express it in simple theorem @hartnn

OpenStudy (briensmarandache):

well is that the same as 1+cot @kryton1212

OpenStudy (anonymous):

cotx = 1/tanx 1+cotx=1+ 1/tanx to continue my working steps, it equal to 1+1/tanx so it's the same

OpenStudy (anonymous):

they are the same*

hartnn (hartnn):

arcsin x + arccos x = pi/2

OpenStudy (anonymous):

o.0 i haven't learnt this haha you help him then :P

hartnn (hartnn):

oops sorry cos(cos^-1(x) + sin^-1(x)) =cos (pi/2) =0

OpenStudy (briensmarandache):

can you do a proof or explain why the sum of those two are equal to pi/2? @hartnn

hartnn (hartnn):

its a standard relation, but let me still prove it for you :)

OpenStudy (briensmarandache):

sweet thank you!!!

hartnn (hartnn):

i got the proof here refer this, http://mathforum.org/library/drmath/view/54161.html if you have any doubts, do ask!